OpenStudy (anonymous):

factoring polynomials completely 75-27y^2

OpenStudy (anonymous):

3(5+3y)(5-3y)

OpenStudy (anonymous):

75-27y^2=3(25-9y^2)=3{5^2-(3y)^2}=3(5+3y)(5-3y) clear?